Assume that exchange rates are the following: Dollar/Pound: 1.60, Yen/Dollar = 100, Yen/Pound = 160

Now assume the following inflation rates over the next year will be: U.S.= 3%, Japan = 1%, and U.K. = 1.5%. What will be the exchange rates one year from now, assuming that Purchasing Power Parity holds:

a) Dollar/Pound b) Yen/Dollar
c) Yen/Pound
Answer #1

If the purchasing power parity holds, US dollar, Japan Yen and UK Pound depriciates by 3%, 1% and 1.5% respectively in absolute.

So, the Dollar/Pound rate would be 1.6x(US inflation)/(UK inflation)=1.6x1.03/1.015=1.623645

So, the Yen/Dollar rate would be 1.6x(Japan inflation)/(US inflation)=100x1.01/1.03=98.0582524

So, the Yen/Pound rate would be 1.6x(Japan inflation)/(UK inflation)=160x1.01/1.015=159.21182266